# Assessment of Spatio-Temporal Predictors in the Presence of Missing and Heterogeneous Data

### v2

@@ Abstract
Deep learning approaches achieve outstanding predictive performance in modeling modern data, despite the increasing complexity and scale. However, evaluating the quality of predictive models becomes more challenging, as traditional statistical assumptions often no longer hold. In particular, spatio-temporal data exhibit dependencies across both time and space, often involving nonlinear dynamics, non-stationarities, and missing observations. As a result, advanced predictors such as spatio-temporal graph neural networks require novel evaluation methodologies. This paper introduces a residual correlation analysis framework designed to assess the optimality of spatio-temporal predictive neural models, particularly in scenarios with incomplete and heterogeneous data. By leveraging the principle that residual correlation indicates information not captured by the model, this framework serves as a powerful tool to identify and localize regions in space and time where model performance can be improved. A key advantage of the proposed approach is its ability to operate under minimal assumptions, enabling robust evaluation of deep learning models applied to multivariate time series, even in the presence of missing and heterogeneous data. The methodology employs tailored spatio-temporal graphs to encode sparse spatial and temporal dependencies within the data and utilizes asymptotically distribution-free summary statistics to pinpoint time intervals and spatial regions where the model underperforms. The effectiveness of the proposed residual analysis is demonstrated through validation on both synthetic and real-world scenarios involving state-of-the-art predictive models.
@@ 1 Introduction
Spatio-temporal predictive neural models fit data by leveraging the inductive biases inherent in relational information derived from the temporal and spatial dimensions of observations (Seo et al., 2018; Bai et al., 2020; Yu et al., 2018; Wu et al., 2019; Pal et al., 2021; Ruiz et al., 2020; Jin et al., 2023). The spatial domain is often represented as a graph, corresponding to structures such as a pixel grid, a 3D mesh, a road map, or a brain network (Shuman et al., 2013; Li et al., 2018; Stankovic et al., 2020). However, the term spatial should be interpreted more broadly, encompassing any functional dependencies among sensors, extending beyond correlations tied to their physical positions (Cini et al., 2023a). Modern spatio-temporal data, such as multivariate time series generated by sensor networks, present significant challenges, including irregular sampling, substantial missing observations, and the dynamic nature of heterogeneous sensors that can be added or removed over time (Alippi, 2014; Montero-Manso and Hyndman, 2021). In such contexts, predictive deep neural models may exhibit varying performance across different regions of the spatio-temporal domain, complicating the analysis of model quality and the detection of unexpected behaviors. Assessing the optimality of these predictors, particularly when leveraging advanced processing neural architectures, remains an intricate task for which no robust and effective methods currently exist.
@@
The quality of deep predictive models is typically assessed using task-specific accuracy metrics, with the squared error being a common choice evaluating the 2-norm of the prediction residuals – the difference between the observed values and the model’s predictions; alternative metrics include absolute (MAE) and relative errors (MAPE) (Gneiting, 2011). These approaches are practical and straightforward, which likely explains their widespread adoption. However, they are inherently comparative, selecting the best model based on statistically superior performance relative to others. Consequently, they provide no direct insight into a model’s optimality or specific guidance on areas that require improvement.
@@
An alternative approach to assessing model quality focuses on analyzing the correlations among prediction residuals rather than their magnitude. The rationale is that correlated residuals indicate structural information not captured by the model (Li, 2003), hence suggesting room for improvement. Over the years, various hypothesis tests have been developed to detect residual dependencies (Durbin and Watson, 1950; Ljung and Box, 1978; Box et al., 2015). Commonly known as randomness or whiteness tests, these tests assess whether residuals exhibit white noise behavior, meaning they lack correlations. However, they rely on strict assumptions, including fully available multivariate time series data, synchronous sampling, and identically distributed data. In practice, though, incomplete and heterogeneous data are the norm, posing significant challenges for existing methods and highlighting the need for more robust tests in such conditions (Zambon and Alippi, 2022).
@@ 1.1 Contributions
This paper presents a novel residual analysis framework for assessing the quality of models designed for spatio-temporal prediction tasks, particularly in scenarios with missing and heterogeneous data, and dynamic (spatial) relational structures. The framework not only detects residual correlations but also pinpoints specific regions where models fail to capture the underlying data-generating process, providing deeper insights beyond traditional evaluation methods. In particular, the paper addresses three key questions about model optimality, intended here as the absence of autocorrelations and cross-dependencies in prediction residuals.
@@
    Is the non-linear neural model optimal in terms of the absence of autocorrelations and cross-dependencies among prediction residuals?
@@
These questions are addressed by partitioning spatial and temporal regions of residuals into tailored subgraphs, followed by the computation of summary statistics, or scores, to quantify potential correlations. These scores measure spatio-temporal correlations by leveraging the AZ-whiteness test (Zambon and Alippi, 2022), which requires neither prior knowledge of the data distribution nor identical distributions across time series.
The key contributions of this paper are as follows:
  • A correlation scoring function that improves over the one introduced in (Zambon and Alippi, 2022), yielding scores that facilitate comparative analysis and interpretation.
@@
The AZ-analysis is validated using synthetic data, and its practical relevance is demonstrated in the real-world scenarios of traffic flow and energy production neural forecasting.
@@ 1.2 Significance and impact
Operating under minimal requirements, the proposed approach is versatile and broadly applicable to real-world scenarios involving deep and graph-based predictors. Notably, it does not impose assumptions on the distribution of residuals or require data to be identically distributed. The only prerequisite is that residuals are centered around zero – an assumption that is typically satisfied in most practical settings.
A key strength of the proposed analysis lies in its graph-based processing, which focuses on residual pairs that are more likely to exhibit correlations. This targeted approach enhances the statistical power of the method, making it more effective in detecting model shortcomings.
While the proposed residual analysis does not directly quantify the extent of potential model improvement, it complements traditional comparative evaluations based on prediction accuracy by providing an independent, metric-agnostic assessment of model quality. Experimental results reveal valuable insights from residual correlation analysis that are not captured through standard prediction error evaluations.
The paper is structured as follows. Section 2 reviews related work. Section 3 introduces the edge scores, the foundation of the proposed residual analysis, and shows how to design the spatio-temporal graph – a multiplex graph that links residual observations based on spatial and temporal relationships. Section 4 provides an overview of the AZ-whiteness test. The paper’s novel contributions are detailed in Section 5 and validated using synthetic data in Section 6. Sections 7 and 8 showcase two real-world applications. Finally, conclusions are presented in Section 9.
@@ 2 Related work
Statistical tests designed to discover correlations in temporal data date back to Durbin and Watson’s investigations (Durbin and Watson, 1950; Durbin, 1970) with variants inspecting longer-range correlations (Box and Pierce, 1970; Ljung and Box, 1978) and multivariate time series (Hosking, 1981; Li et al., 2019).
More general statistical tests leverage graph-based structures to connect data points. One of the earliest examples is the two-sample Friedman-Rafsky test (Friedman and Rafsky, 1979), which assesses the equality of distributions. Building on this foundation, subsequent research has extended these methods to change-point detection tests and statistical tests applicable to high-dimensional and non-Euclidean data (Chen and Zhang, 2013; Chu and Chen, 2019; Chen and Chu, 2023).
Following the approach in (Wald and Wolfowitz, 1940; Geary, 1970), which employs sign statistics to design non-parametric tests, Zambon and Alippi (2022) introduced a whiteness test to evaluate the overall presence of spatial and/or temporal correlation in prediction residuals.
To the best of our knowledge, this paper is the first one providing a comprehensive spatio-temporal analysis of correlation patterns at the node (time series), temporal, and local spatio-temporal levels.
@@ 3.1 Spatio-temporal predictions
where $\boldsymbol{\mathbf{x}}_{t,v}\in\mathbb{R}^{d_{x}}$ is a stochastic vector associated with observation of the $v$-th time series at time step $t$. We may refer to $v\in V_{t}$ as the sensor associated with the $v$-th time series from the sensor set $V_{t}$ available at time step $t$; denote $\boldsymbol{\mathbf{x}}_{t}\doteq\{\boldsymbol{\mathbf{x}}_{t,v}:v\in V_{t}\}$ and $V\doteq\bigcup_{t}V_{t}$. Consider a family of predictive models $f_{\theta}$ with parameter vector $\theta$ trained to predict at every time step $t$ target $\boldsymbol{\mathbf{y}}_{t,v}\in\mathbb{R}^{d_{y}}$, for all $v$, starting from a set of inputs in $\boldsymbol{\mathbf{x}}$; exogenous variables, such as those containing information about different types of sensors, are assumed encoded in $\boldsymbol{\mathbf{x}}$. A common example of a predictive task is time series forecasting (Cini et al., 2023a). In this context, $\boldsymbol{\mathbf{y}}_{t}$ corresponds to the window of subsequent observations $\boldsymbol{\mathbf{x}}_{t:t+H}$, therefore $d_{y}=d_{x}\times H$; notation $\boldsymbol{\mathbf{x}}_{a:b}$ indicates $\{\boldsymbol{\mathbf{x}}_{t}:a\leq t<b\}$. The presented formulation naturally handles missing data in $\boldsymbol{\mathbf{x}}$ using the sets $V_{t}$, which may not necessarily be equal to $V$. Moreover, no specific homogeneity assumption is made on the data, and the capability to support heterogeneous sensors will become evident from the discussion in Section 4.
In this paper, we consider the existence of functional dependencies among observations either available or extracted from the data itself (Cini et al., 2023c). We embed dependencies at time $t$ in graph
@@
where node set $V_{t}$, edge topology $E_{t}$ and edge weights $\boldsymbol{\mathbf{w}}_{t}$ can change over time. A visual representation of the set of time series $\boldsymbol{\mathbf{x}}$ and their relations in $\{\boldsymbol{\mathbf{g}}_{t}\}$ is provided in Figure LABEL:fig:st-data.
Model predictions $\hat{\boldsymbol{\mathbf{y}}}_{t}=f_{\hat{\theta}}(\boldsymbol{\mathbf{x}}_{t-W:t})$ are typically obtained from a sliding window of input data, defined as $\boldsymbol{\mathbf{x}}_{t-W:t}$, encompassing the past $W$ time steps. The parameter vector $\hat{\theta}$ is learned by minimizing a prediction loss function, such as mean squared error (MSE) or mean absolute error (MAE), between the predicted values $\hat{\boldsymbol{\mathbf{y}}}_{t,v}$ and the ground truth targets $\boldsymbol{\mathbf{y}}_{t,v}$. More broadly, predictions at time step $t$ can account for topological information: $\hat{\boldsymbol{\mathbf{y}}}_{t}=f_{\hat{\theta}}(\boldsymbol{\mathbf{g}}_{t-W:t})$, as done in graph-based models like spatio-temporal graph neural networks (STGNNs) (Cini et al., 2023a), as well as past predictions $\hat{\boldsymbol{\mathbf{y}}}_{t-W:t}$ and target values $\boldsymbol{\mathbf{y}}_{t-W:t}$. Define prediction residuals as
@@
for all $v$ and $t$; the $i$-th component $\boldsymbol{\mathbf{r}}_{t,v}(i)$ of vector $\boldsymbol{\mathbf{r}}_{t,v}$ is the residual associated with the $i$-th target variable $\boldsymbol{\mathbf{y}}_{t,v}(i)$.
@@ 3.2 Spatio-temporal graphs
obtained by stacking all graphs $\boldsymbol{\mathbf{g}}_{t}$, for all $t=1,2,\dots,T$, and joining each node at time $t$ with itself at time $t-1$ as shown in Figure 2; in the figure, solid gray lines denote edges in $\boldsymbol{\mathbf{g}}_{t}$, while dashed gray lines indicate temporal edges across consecutive graphs. Moreover, $\boldsymbol{\mathbf{g}}^{*}$ is augmented with the residuals $\boldsymbol{\mathbf{r}}$ as node signals and a set $\boldsymbol{\mathbf{w}}$ of edge weights. As we will see in the sequel, the spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ provides a convenient static representation of the observed system dynamics.
@@
$E_{\rm sp}$ collects all (spatial) edges in sets $E_{t}$, for all $t$, regardless of their orientation whenever it applies, while $E_{\rm tm}$ is the set of edges connecting corresponding nodes along the temporal dimension. Note that edge directions and self-loops are excluded in (9). However, this have no impact on the following analysis, as the statistics derived from equation (16) disregard directional information. Secondly, our focus is on the correlations between different observations in both time and space. Finally, if $\boldsymbol{\mathbf{g}}_{t}$ is undirected, weight $w_{u_{t},v_{t}}$ of spatial edge $\{u_{t},v_{t}\}\in E_{\rm sp}$ in $\boldsymbol{\mathbf{g}}^{*}$ is set equal to weight $w_{t,\{u,v\}}$ of $\boldsymbol{\mathbf{g}}_{t}$, i.e., the weight corresponding to edge $\{u,v\}\in E_{t}$. Conversely, for a directed graph $\boldsymbol{\mathbf{g}}_{t}$, $w_{u_{t},v_{t}}$ equals either $w_{t,(u,v)}$ or $w_{t,(v,u)}$, depending on which directional edge is in $E_{t}$, and $w_{u_{t},v_{t}}=w_{t,(u,v)}+w_{t,(v,u)}$ if both $(u,v)$ and $(v,u)$ are present. The weights for temporal edges can be defined arbitrarily or, e.g., to balance the overall impact of the spatial and temporal edges, as discussed in Section 4.

### v3

# arXiv:2302.01701v3[stat.ML] 27 Feb 2026

@@ Abstract
Deep learning methods achieve remarkable predictive performance in modeling complex, large-scale data. However, assessing the quality of derived models has become increasingly challenging, as more classical statistical assumptions may no longer apply. These difficulties are particularly pronounced for spatio-temporal data, which exhibit dependencies across both space and time and are often characterized by nonlinear dynamics, time variance, and missing observations, hence calling for new accuracy assessment methodologies. This paper introduces a residual correlation analysis framework for assessing the optimality of spatio-temporal relational-enabled neural predictive models, notably in settings with incomplete and heterogeneous data. By leveraging the principle that residual correlation indicates information not captured by the model, enabling the identification and localization of regions in space and time where predictive performance can be improved. A strength of the proposed approach is that it operates under minimal assumptions, allowing also for robust evaluation of deep learning models applied to multivariate time series, even in the presence of missing and heterogeneous data. In detail, the methodology constructs tailored spatio-temporal graphs to encode sparse spatial and temporal dependencies and employs asymptotically distribution-free summary statistics to detect time intervals and spatial regions where the model underperforms. The effectiveness of what proposed is demonstrated through experiments on both synthetic and real-world datasets using state-of-the-art predictive models.
footnotetext:
This article is published in Neurocomputing (Elsevier), 2026, under the CC BY 4.0 license. The final version of record is available at: https://doi.org/10.1016/j.neucom.2026.132963
@@ 1 Introduction
Spatio-temporal predictive neural models fit data by leveraging the inductive biases inherent in possibly latent relational information derived from the temporal and spatial dimensions of observations (Seo et al., 2018; Bai et al., 2020; Yu et al., 2018; Wu et al., 2019; Pal et al., 2021; Ruiz et al., 2020; Jin et al., 2024). The spatial domain is often represented as a graph, capturing structures such as pixel grids, 3D meshes, road maps, or brain networks (Shuman et al., 2013; Li et al., 2018; Stanković et al., 2020). However, the term spatial should be interpreted more broadly, encompassing functional dependencies among sensors that go beyond correlations tied to their physical positions (Cini et al., 2025). Modern spatio-temporal data, such as multivariate time series generated by sensor networks, present significant challenges, including irregular sampling, substantial missing observations, and the time variant nature of heterogeneous sensors that can be added or removed over time (Alippi, 2014; Montero-Manso and Hyndman, 2021). In these contexts, predictive deep neural models may perform inconsistently across different regions of the spatio-temporal domain, complicating model quality assessments and the detection of unexpected behaviors. Assessing the optimality of these predictors, particularly when leveraging advanced processing neural architectures, remains an intricate task for which no robust and effective methods currently exist.
@@
The quality of deep predictive models is typically assessed using task-specific accuracy metrics, with the squared error being a common choice, evaluating the 2-norm of the prediction residuals – the difference between the observed values and the model’s predictions. Alternative metrics include absolute (MAE) and relative errors (MAPE) (Gneiting, 2011). These approaches are practical and straightforward, likely contributing to their widespread adoption. However, they are inherently comparative, selecting the best model based on statistically superior performance relative to others. Consequently, they offer no direct insight into model optimality or specific guidance on areas needing improvement.
@@
An alternative approach to assessing model quality focuses on analyzing the correlations among prediction residuals rather than their magnitude. The rationale is that correlated residuals indicate structural information that the model has failed to capture (Li, 2003), hence suggesting room for improvement. Over the years, various hypothesis tests have been developed to detect residual dependencies (Durbin and Watson, 1950; Ljung and Box, 1978; Box et al., 2015). Commonly known as randomness or whiteness tests, these tests assess whether residuals exhibit white noise behavior, meaning they lack correlations. However, they rely on strict assumptions, including fully available multivariate time series, synchronous sampling, and identically distributed data. In practice, however, incomplete and heterogeneous data are the norm, creating significant challenges for existing methods and underscoring the need for more robust tests (Zambon and Alippi, 2022).
@@ 1.1 Contributions
This paper presents a novel residual analysis framework for assessing the quality of models designed for spatio-temporal prediction tasks, particularly in scenarios with missing and heterogeneous data and dynamic (spatial) relational structures. The framework not only detects residual correlations but also pinpoints specific regions where models fail to capture the underlying data-generating process, offering a more nuanced understanding than traditional approaches. In particular, the paper addresses three key questions about model optimality, intended here as the absence of autocorrelations and cross-dependencies in prediction residuals.
@@
    Is the nonlinear neural model optimal in terms of the absence of autocorrelations and cross-dependencies among prediction residuals?
@@
The proposed residual analysis, referred to as AZ-analysis, addresses these questions by building on the statistic underlying the AZ-whiteness test (Zambon and Alippi, 2022). This enables the method to inherit the flexibility of the original test in handling complex spatio-temporal data, without requiring prior knowledge of the data distribution or identical distributions across time series. At the same time, AZ-analysis is conceptually distinct from the AZ-whiteness test, repurposing the underlying test statistic to construct interpretable, region-wise correlation measures. In particular, AZ-analysis introduces tailored subgraphs that partition residuals into spatial and temporal regions and computes summary statistics, or scores, to quantify local correlation and enable comparisons across different regions of the data. Building on these new elements, the paper develops a comprehensive framework for analyzing residual correlations at multiple levels, enabling the identification and localization of temporal drifts, node-specific dynamics not captured by the model, and anomalies rooted in data acquisition. The AZ-analysis also offers practical guidelines to interpret the results and address the identified correlation patterns. The key contributions of this paper are as follows:
@@
The AZ-analysis is validated on synthetic data, and its practical relevance is demonstrated in real-world scenarios of traffic flow and energy production forecasting.
@@ 1.2 Significance and impact
Requiring minimal assumptions, the proposed approach is versatile and broadly applicable to real-world scenarios involving deep and graph-based predictors. Notably, it does not impose assumptions on the distribution of residuals or require data to be identically distributed. The only prerequisite is that residuals are centered at zero – an assumption that is typically satisfied in most practical settings. This is one of the main advances with respect to the related literature, as further elaborated in Section 2. Another key strength of the proposed analysis lies in its graph-based processing, which focuses on residual pairs that are more likely to exhibit correlations. This targeted approach enhances the statistical power of the method, making it more effective in detecting model shortcomings.
Although the proposed residual analysis does not quantify the magnitude of potential model improvements, it complements traditional accuracy-based evaluations. As demonstrated in Section 7, it provides an independent, metric-agnostic assessment of model quality. In particular, the experimental results reveal valuable insights from residual correlation analysis that standard prediction error evaluations fail to capture.
The paper is structured as follows. Section 2 reviews related work. Section 3 introduces the edge scores, the foundation of the proposed residual analysis, and shows how to design the spatio-temporal graph – a multiplex graph linking residual observations through spatial and temporal relationships. Section 4 provides an overview of the AZ-whiteness test. The paper’s novel contributions are detailed in Section 5 and validated using synthetic data in Section 6. Sections 7 and 8 showcase two real-world applications. Finally, conclusions are presented in Section 9.
@@ 2 Related work
Statistical tests designed to discover correlations in temporal data date back to Durbin and Watson’s investigations (Durbin and Watson, 1950; Durbin, 1970). Most of the literature has focused on univariate time series (Geary, 1970; Drouiche, 2002), although variants have appeared for inspecting longer-range correlations (Box and Pierce, 1970; Ljung and Box, 1978) and multivariate time series (Durbin, 1957; Hosking, 1981; Li et al., 2019). These (classical) tests, however, often rely on strong assumptions, such as complete data, relatively low dimensionality, and identical distribution of observations, that may not hold in complex modern datasets. More general statistical tests that leverage graph-based structures to connect data points can be considered. One of the earliest examples is the two-sample Friedman-Rafsky test (Friedman and Rafsky, 1979), which assesses the equality of distributions by looking at whether points sharing an edge belong to the same sample. Building on this foundation, subsequent research has extended these methods to change-point detection tests and statistical tests applicable to high-dimensional and non-Euclidean data (Chen and Zhang, 2013; Chu and Chen, 2019; Chen and Chu, 2023). By exploiting general data relationships, these graph-based tests remain effective in settings where traditional parametric tests break down, enabling hypothesis testing on complex data objects without requiring stringent assumptions about data dimensionality.
Orthogonal to the above, non-parametric tests using runs and sign statistics have long been used to check randomness in a sequence. The Wald and Wolfowitz (1940) runs test, and later the Friedman and Rafsky (1979) test, introduced the idea of counting how many adjacent observations come from the same sample. Geary (1970) employed sign statistics on prediction residuals as an indicator of serial correlation. More recently, Zambon and Alippi (2022) generalized the idea to spatio-temporal data, introducing a whiteness test to evaluate the overall presence of both spatial and temporal correlation in prediction residuals.
However, classical hypothesis tests provide only global summary statistics and do not localize the sources of correlation within space or time. With the recent rise of spatio-temporal prediction models in deep learning (Jin et al., 2024; Cini et al., 2025), it has become necessary – and is the focus of this paper – to assess model quality at a finer-grained level, to better understand model behavior amid such complexity. The works of Papadimitriou et al. (2006) and Zhao (2015) are two examples that have addressed local correlation patterns in the temporal domain. In contrast, the local indicators of spatial association (LISA) framework (Anselin, 1995) is a robust approach designed for spatial data, exploring local autocorrelation through statistics such as Moran’s I (Moran, 1950), Geary’s C (Geary, 1954), and multivariate ones (Anselin, 2019). However, these approaches rarely consider scenarios involving missing or heterogeneous data distributions.
To the best of our knowledge, this paper is the first to provide a comprehensive spatio-temporal analysis of correlation patterns at the node (time series), temporal, and local spatio-temporal levels. In particular, the proposed AZ-analysis combines the local insight offered by the LISA framework with the flexibility of sign-based statistics, enabling wide applicability under mild assumptions.
@@ 3.1 Spatio-temporal predictions
where $\boldsymbol{\mathbf{x}}_{t,v}\in\mathbb{R}^{d_{x}}$ is a stochastic vector representing the observation of the $v$-th time series at time step $t$. We may refer to $v\in V_{t}$ as the multi-channel sensor associated with the $v$-th time series from the sensor set $V_{t}$ available at time step $t$; denote $\boldsymbol{\mathbf{x}}_{t}\doteq\{\boldsymbol{\mathbf{x}}_{t,v}:v\in V_{t}\}$ and $V\doteq\bigcup_{t}V_{t}$. Consider a family of predictive models $f_{\theta}$ with parameter vector $\theta$ trained to predict, at every time step $t$, target $\boldsymbol{\mathbf{y}}_{t,v}\in\mathbb{R}^{d_{y}}$ for all $v$. The inputs to these models come from $\boldsymbol{\mathbf{x}}$. Exogenous variables, such as those describing different types of sensors, are assumed to be encoded in $\boldsymbol{\mathbf{x}}$. A common example of a predictive task is time series forecasting (Cini et al., 2025): in this context, $\boldsymbol{\mathbf{y}}_{t}$ corresponds to the window of subsequent observations $\boldsymbol{\mathbf{x}}_{t:t+H}$, therefore $d_{y}=d_{x}\times H$; notation $\boldsymbol{\mathbf{x}}_{a:b}$ indicates $\{\boldsymbol{\mathbf{x}}_{t}:a\leq t<b\}$. The presented formulation naturally handles missing data in $\boldsymbol{\mathbf{x}}$ using the sets $V_{t}$, which may not necessarily be equal to $V$. Moreover, no specific homogeneity assumptions are imposed on the data, and the ability to handle heterogeneous sensors will be evident from the discussion in Section 4.
In this paper, we consider the existence of functional dependencies among observations either available or extracted from the data itself (Cini et al., 2023b). We encode these dependencies at time $t$ in graph
@@
where node set $V_{t}$, edge topology $E_{t}$ and edge weights $\boldsymbol{\mathbf{w}}_{t}$ can change over time. A visual representation of the set of time series $\boldsymbol{\mathbf{x}}$ and their relations in $\{\boldsymbol{\mathbf{g}}_{t}\}$ is provided in Figure 1.
Model predictions $\hat{\boldsymbol{\mathbf{y}}}_{t}=f_{\hat{\theta}}(\boldsymbol{\mathbf{x}}_{t-W:t})$ are typically obtained from a sliding window of input data, defined as $\boldsymbol{\mathbf{x}}_{t-W:t}$, encompassing the past $W$ time steps. The parameter vector $\hat{\theta}$ is learned by minimizing a prediction loss function, such as mean squared error (MSE) or mean absolute error (MAE), between the predicted values $\hat{\boldsymbol{\mathbf{y}}}_{t,v}$ and the ground truth targets $\boldsymbol{\mathbf{y}}_{t,v}$. More broadly, predictions at time step $t$ can incorporate topological information ($\hat{\boldsymbol{\mathbf{y}}}_{t}=f_{\hat{\theta}}(\boldsymbol{\mathbf{g}}_{t-W:t})$) as in graph-based models like spatio-temporal graph neural networks (STGNNs) (Cini et al., 2025). Predictions may also depend on past predictions $\hat{\boldsymbol{\mathbf{y}}}_{t-W:t}$ and previous target values $\boldsymbol{\mathbf{y}}_{t-W:t}$. Define prediction residuals as
@@
for all $v$ and $t$. The $i$-th component, $\boldsymbol{\mathbf{r}}_{t,v}(i)$, of vector $\boldsymbol{\mathbf{r}}_{t,v}$ is the residual associated with the $i$-th target variable $\boldsymbol{\mathbf{y}}_{t,v}(i)$.
@@ 3.2 Spatio-temporal graphs
constructed by stacking all graphs $\boldsymbol{\mathbf{g}}_{t}$, for all $t=1,2,\dots,T$. Each node at time $t$ is connected to itself at time t1t-1, as shown in Figure 2. In the figure, solid gray lines denote edges in $\boldsymbol{\mathbf{g}}_{t}$, while dashed gray lines indicate temporal edges across consecutive graphs. Moreover, $\boldsymbol{\mathbf{g}}^{*}$ is augmented with the residuals $\boldsymbol{\mathbf{r}}$ as node signals and a set $\boldsymbol{\mathbf{w}}$ of edge weights. As we will see later, the spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ offers a convenient static representation of the observed system dynamics.
@@
$E_{\rm sp}$ collects all (spatial) edges in sets $E_{t}$, for all $t$, regardless of their orientation whenever it applies, while $E_{\rm tm}$ is the set of edges connecting corresponding nodes along the temporal dimension. Note that edge directions and self-loops are excluded in (9). However, this has no impact on the following analysis, as the statistics derived from equation (16) disregard directional information. Secondly, our focus is on the correlations between different observations in both time and space. Finally, if $\boldsymbol{\mathbf{g}}_{t}$ is undirected, weight $w_{u_{t},v_{t}}$ of spatial edge $\{u_{t},v_{t}\}\in E_{\rm sp}$ in $\boldsymbol{\mathbf{g}}^{*}$ is set equal to weight $w_{t,\{u,v\}}$ of $\boldsymbol{\mathbf{g}}_{t}$, i.e., the weight corresponding to edge $\{u,v\}\in E_{t}$. Conversely, for a directed graph $\boldsymbol{\mathbf{g}}_{t}$, the weight $w_{u_{t},v_{t}}$ equals $w_{t,(u,v)}$ or $w_{t,(v,u)}$, depending on which directional edge is in $E_{t}$. If both $(u,v)$ and $(v,u)$ are present, then $w_{u_{t},v_{t}}=w_{t,(u,v)}+w_{t,(v,u)}$. The weights for temporal edges can be defined arbitrarily or, e.g., to balance the overall impact of the spatial and temporal edges, as discussed in Section 4.
For clarity of presentation, Table 1 summarizes the main notation adopted throughout the paper.
Table 1: Summary of notation used in the paper.
Symbol Description
$\boldsymbol{\mathbf{x}}_{t,v}\in\mathbb{R}^{d_{x}}$ Observation of sensor/node $v$ at time $t$
$\boldsymbol{\mathbf{y}}_{t,v},\ \hat{\boldsymbol{\mathbf{y}}}_{t,v}\in\mathbb{R}^{d_{y}}$ Target and model prediction at $(t,v)$
$\boldsymbol{\mathbf{r}}_{t,v}\doteq\boldsymbol{\mathbf{y}}_{t,v}-\hat{\boldsymbol{\mathbf{y}}}_{t,v}$ Prediction residual at $(t,v)$
$\boldsymbol{\mathbf{r}}\doteq\{\boldsymbol{\mathbf{r}}_{t,v}:\forall v,t\}$ Collection of all prediction residuals
$\boldsymbol{\mathbf{g}}_{t},\ V_{t},\ E_{t}$ Graph at time $t$ and corresponding node and edge sets
$\boldsymbol{\mathbf{w}}_{t},\ w_{t,e}$ Set of edge weights of $\boldsymbol{\mathbf{g}}_{t}$ and weight of edge $e\in E_{t}$
$\boldsymbol{\mathbf{g}}^{*}\doteq(V^{*},E^{*},\boldsymbol{\mathbf{w}},\boldsymbol{\mathbf{r}})$ Multiplex graph with $\boldsymbol{\mathbf{r}}$ as node attributes
$E_{\mathrm{sp}},\ E_{\mathrm{tm}}$ Spatial and temporal edge sets of $\boldsymbol{\mathbf{g}}^{*}$
$\boldsymbol{\mathbf{s}}=(V_{s},E_{s},\boldsymbol{\mathbf{w}}_{s},\boldsymbol{\mathbf{r}}_{s})$ Generic subgraph of $\boldsymbol{\mathbf{g}}^{*}$
$|E_{s}|,\ \lVert\boldsymbol{\mathbf{w}}_{s}\rVert_{1},\ \lVert\boldsymbol{\mathbf{w}}_{s}\rVert_{2}$ Number of edges, 1- and 2-norm of the edge weights of $\boldsymbol{\mathbf{s}}$
$C_{\lambda}(\boldsymbol{\mathbf{s}}),\ c_{\lambda}(\boldsymbol{\mathbf{s}})$ AZ-whiteness test statistic and correlation score on $\boldsymbol{\mathbf{s}}$
$c_{\lambda}(v),\ c_{\lambda}(t),\ c_{\lambda}(v,t)$ Node, time, and local scores at node $v$ and time $t$
$\lambda\in[0,1]$ Parameter trading off spatial and temporal contributions
$\gamma>0,\ \alpha\in(0,1)$ Threshold and significance level of the AZ-whiteness test